OpenStudy (anonymous):

what is the derivative of 60+15cos(4pi*t)

zepdrix (zepdrix):

\[\huge f(t)=60+15\cos(4\pi t)\] \[\huge f'(t)=0+15\left(-\sin(4\pi t)\right)(4\pi t)'\]Apply the chain rule, you make a copy of the inside, takes it's derivative and multiply it on the outside. The prime is to show that we still need to take it's derivative.